# HyperGraphRAG: Retrieval-Augmented Generation via Hypergraph-Structured Knowledge Representation

### v2

@@ Abstract
Standard Retrieval-Augmented Generation (RAG) relies on chunk-based retrieval, whereas GraphRAG advances this approach by graph-based knowledge representation. However, existing graph-based RAG approaches are constrained by binary relations, as each edge in an ordinary graph connects only two entities, limiting their ability to represent the n-ary relations ($n\geq 2$) in real-world knowledge. In this work, we propose HyperGraphRAG, a novel hypergraph-based RAG method that represents n-ary relational facts via hyperedges, and consists of knowledge hypergraph construction, retrieval, and generation. Experiments across medicine, agriculture, computer science, and law demonstrate that HyperGraphRAG outperforms both standard RAG and previous graph-based RAG methods in answer accuracy, retrieval efficiency, and generation quality. Our data and code are publicly available at https://github.com/LHRLAB/HyperGraphRAG.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) [10, 6] has advanced knowledge-intensive tasks by integrating knowledge retrieval with large language models (LLMs) [17, 28], thereby enhancing factual awareness and generation accuracy. Standard RAG typically relies on chunk-based retrieval, segmenting documents into fixed-length text chunks retrieved via dense vector similarity, which overlooks the relationships between entities. Recently, GraphRAG [2] has emerged as a promising direction that structures knowledge as a graph to capture inter-entity relations, with the potential to improve retrieval efficiency and knowledge-driven generation [18].
However, since each edge in an ordinary graph connects only two entities, existing graph-based RAG approaches [2, 7, 1, 8] are all restricted to binary relations, making them insufficient for modeling the n-ary relations among more than two entities that are widespread in real-world domain knowledge [25]. For example, in the medical domain, as illustrated in Figure 2, representing the fact that “Male hypertensive patients with serum creatinine levels between 115–133 $\mu$mol/L are diagnosed with mild serum creatinine elevation” requires decomposing it into several binary relational triples, such as Gender:(Hypertensive patient, Male) and Diagnosed_with:(Hypertensive patient, Mild serum creatinine elevation), leading to representation loss and sparsity during the conversion.
@@
To address these limitations, we propose HyperGraphRAG, as illustrated in Figure 1, a novel graph-based RAG method built upon hypergraph-structured knowledge representation. In contrast to prior graph-based RAG methods constrained to binary relations, HyperGraphRAG leverages hyperedges to represent n-ary relational facts, where each hyperedge connects $n$ entities ($n\geq 2$), e.g. Hyperedge:(Hypertensive patient, Male, Serum creatinine levels between 115–133 $\mu$mol/L, Mild serum creatinine elevation), and each hyperedge is expressed through natural language descriptions. This design ensures knowledge completeness, structural expressiveness, and inferential capability, thereby providing more comprehensive support for knowledge-intensive applications.
@@
To validate the effectiveness, we conduct experiments in multiple knowledge-intensive domains [7], including medicine, agriculture, computer science, and law. Results demonstrate that HyperGraphRAG outperforms standard RAG and previous graph-based RAG methods in answer accuracy, retrieval efficiency, and generation quality, showcasing its strong potential for real-world applications.
@@ 2 Related Work
Graph-based RAG. GraphRAG [2] is the first graph-based RAG method that improves LLM generation via graph-based retrieval. Based on GraphRAG, several methods [26, 22, 11, 4, 23] focus on building graph-based RAG for different applications. LightRAG [7] enhances efficiency via graph indexing and updates. PathRAG [1] and HippoRAG2 [8] refine retrieval with path pruning and Personalized PageRank. However, all rely on binary relations, limiting knowledge expressiveness. In this work, we propose HyperGraphRAG, the first graph-based RAG method via hypergraph-structured knowledge representation. We compare several existing methods with HyperGraphRAG in Table 1.
Hypergraph Representation. Hypergraph-structured knowledge representation aims to overcome ordinary graph’s limitations in modeling n-ary relations [15]. Early methods [25, 27, 12, 21] employ various embedding techniques to represent n-ary relational entities. Later methods [5, 24, 14] utilize GNN or attention to enhance embedding. However, existing methods mainly focus on link prediction, while hypergraphs also show potential for enhancing knowledge representation in graph-based RAG.
@@
GraphRAG [2] $\mathcal{K}=S=\{s_{g}\mid g\in\text{Community}(G)\}$,
@@
LightRAG [7] $\mathcal{K}=G=(V,E)$,
where $V$ $E$ are entity & relation sets.
@@
HippoRAG2 [8] $\mathcal{K}=G=(V\cup M,\;E)$,
where $V$ $M$ are phrase & passage nodes.
@@ 3 Preliminaries
Definition 3: Hypergraph. A hypergraph $G_{H}=(V,E_{H})$ [29] is a generalized graph, where $V$ is the entity set, $E_{H}$ is the hyperedge set, and each hyperedge $e_{H}\in E_{H}$ connects 2 or more entities:
@@ 4.1 Knowledge Hypergraph Construction
N-ary Relation Extraction. To construct the knowledge hypergraph $G_{H}$, our first step is to extract multiple n-ary relational facts $F_{n}$ from natural language documents $d\in K$. Unlike traditional hyper-relations [21], events [13], or other n-ary relation models [15], in the era of LLMs, to preserve richer and more diverse n-ary relations among entities, we propose a new n-ary relation representation $F_{n}=(e_{H},V_{e_{H}})$, utilizing natural language descriptions, instead of structured relations, to represent hyperedges $e_{H}$ among multiple entities $V_{e_{H}}$ as follows.
@@ 5.1 Experimental Setup
Datasets. To evaluate the performance of HyperGraphRAG across multiple domains, we select four knowledge contexts from UltraDomain [19], as used in LightRAG [7]: Agriculture, Computer Science (CS), Legal, and a mixed domain (Mix). In addition, we include the latest international hypertension guidelines [16] as the foundational knowledge for the Medicine domain. For each of the five domains, we sample knowledge fragments one, two, and three hops away to construct questions with ground-truth answers verified by human annotators. We then categorize the questions into Binary Source and N-ary Source, based on whether the sampled knowledge of the question contains facts among $n$ entities ($n>2$). More details can be found in Appendix D.
Baselines. We compare HyperGraphRAG against six publicly available baseline methods: NaiveGeneration [17], which directly generates responses using LLM; StandardRAG [6], a traditional chunk-based RAG approach; GraphRAG [2], LightRAG [7], PathRAG [1], and HippoRAG2 [8], which are four selected available graph-based RAG methods described in Table 1. To ensure fairness, we use the same generation prompt, which can be found in Appendix E.
Evaluation Metrics. We evaluate the answer accuracy, retrieval efficiency, and generation quality of HyperGraphRAG and its baselines using 3 key metrics: F1, Retrieval Similarity (R-S), and Generation Evaluation (G-E). F1 measures word-level similarity between the generated answer and the ground-truth answer, following FlashRAG [9]. R-S assesses the semantic similarity between the retrieved knowledge and the ground-truth knowledge used to construct the question, in line with RAGAS [3]. G-E, inspired by HelloBench [20], is a metric that uses LLM-as-a-judge to evaluate generation quality in 7 dimensions and reports the average score. Details are provided in Appendix E.
Implementation Details. We use OpenAI’s GPT-4o-mini for extraction and generation, and text-embedding-3-small for vector. During retrieval, we set the following parameters: entity retrieval $k_{V}=60$, $\tau_{V}=50$; hyperedge retrieval $k_{H}=60$, $\tau_{H}=5$; and chunk retrieval $k_{C}=6$, $\tau_{C}=0.5$. All experiments were conducted on a server with an 80-core CPU and 512GB RAM.
@@ 5.5 Analysis of Hypergraph Retrieval Efficiency (RQ4)
5.6 Analysis of Hypergraph-Guided Generation Equality (RQ5)
@@ B.3 Proof of Proposition 3
[…]
@@
[…]
@@ Appendix E Evaluation Details
[…]
@@
[…]
@@
[…]
@@ Appendix G Hyperparameter Settings
For all methods, we adopt a unified set of hyperparameters for all models across both the main evaluation in Table 2 and the time/cost experiments in Table 3 to ensure fair and consistent comparison. For chunk-based methods (e.g., StandardRAG), we retrieve the top-6 chunks using dense similarity. For graph-based methods, including GraphRAG, LightRAG, PathRAG, and HippoRAG2, we retrieve the top-60 relevant elements according to their respective retrieval strategies. HyperGraphRAG performs dual top-60 retrieval over entities and hyperedges, followed by neighborhood expansion. All methods are run using 16 parallel cores and the same generation model (GPT-4o-mini) with temperature 1.0 and a maximum generation length of 32k tokens. Table 4 summarizes the detailed hyperparameter configurations used throughout our experiments.
@@
StandardRAG Chunk 6 chunks 16 GPT-4o-mini

### v3

@@ Abstract
Standard Retrieval-Augmented Generation (RAG) relies on chunk-based retrieval, whereas GraphRAG advances this approach by graph-based knowledge representation. However, existing graph-based RAG approaches are constrained by binary relations, as each edge in an ordinary graph connects only two entities, limiting their ability to represent the n-ary relations ($n\geq 2$) in real-world knowledge. In this work, we propose HyperGraphRAG, the first hypergraph-based RAG method that represents n-ary relational facts via hyperedges. HyperGraphRAG consists of a comprehensive pipeline, including knowledge hypergraph construction, retrieval, and generation. Experiments across medicine, agriculture, computer science, and law demonstrate that HyperGraphRAG outperforms both standard RAG and previous graph-based RAG methods in answer accuracy, retrieval efficiency, and generation quality. Our data and code are publicly available111 https://github.com/LHRLAB/HyperGraphRAG.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) (RAG, ; RAGsurvey, ) has advanced knowledge-intensive tasks by integrating knowledge retrieval with large language models (LLMs) (GPT4, ; LLMsurvey, ), thereby enhancing factual awareness and generation accuracy. Standard RAG typically relies on chunk-based retrieval, segmenting documents into fixed-length text chunks retrieved via dense vector similarity, which overlooks the relationships between entities. Recently, GraphRAG (GraphRAG, ) has emerged as a promising direction that structures knowledge as a graph to capture inter-entity relations, with the potential to improve retrieval efficiency and knowledge-driven generation (KGLLMsurvey, ).
[…]
@@
[…]
@@
To validate the effectiveness, we conduct experiments in multiple knowledge-intensive domains (LightRAG, ), including medicine, agriculture, computer science, and law. Results demonstrate that HyperGraphRAG outperforms standard RAG and previous graph-based RAG methods in answer accuracy, retrieval efficiency, and generation quality, showcasing its strong potential for real-world applications.
@@ 2 Related Work
Graph-based RAG. GraphRAG (GraphRAG, ) is the first graph-based RAG method that improves LLM generation via graph-based retrieval. Based on GraphRAG, several methods (MedGraphRAG, ; OG-RAG, ; KAG, ; MiniRAG, ; PIKE-RAG, ) focus on building graph-based RAG for different applications. LightRAG (LightRAG, ) enhances efficiency via graph indexing and updates. PathRAG (PathRAG, ) and HippoRAG2 (HippoRAG2, ) refine retrieval with path pruning and Personalized PageRank. However, all rely on binary relations, limiting knowledge expressiveness. In this work, we propose HyperGraphRAG, the first graph-based RAG method via hypergraph-structured knowledge representation. We compare several existing methods with HyperGraphRAG in Table 1.
Hypergraph Representation. Hypergraph-structured knowledge representation aims to overcome ordinary graph’s limitations in modeling n-ary relations (Text2NKG, ). Early methods (m-TransH, ; RAE, ; n-TuckER, ; HINGE, ) employ various embedding techniques to represent n-ary relational entities. Later methods (StarE, ; GRAN, ; HAHE, ) utilize GNN or attention to enhance embedding. However, existing methods mainly focus on link prediction, while hypergraphs also show potential for enhancing knowledge representation in graph-based RAG.
@@
GraphRAG (GraphRAG, ) $\mathcal{K}=S=\{s_{g}\mid g\in\text{Community}(G)\}$,
@@
LightRAG (LightRAG, ) $\mathcal{K}=G=(V,E)$,
where $V$ & $E$ are entity & relation sets.
@@
HippoRAG2 (HippoRAG2, ) $\mathcal{K}=G=(V\cup M,\;E)$,
where $V$ & $M$ are phrase & passage nodes.
@@ 3 Preliminaries
Definition 3: Hypergraph. A hypergraph $G_{H}=(V,E_{H})$ (Hypergraph, ) is a generalized graph, where $V$ is the entity set, $E_{H}$ is the hyperedge set, and each hyperedge $e_{H}\in E_{H}$ connects 2 or more entities:
@@ 4.1 Knowledge Hypergraph Construction
N-ary Relation Extraction. To construct the knowledge hypergraph $G_{H}$, our first step is to extract multiple n-ary relational facts $F_{n}$ from natural language documents $d\in K$. Unlike traditional hyper-relations (HINGE, ), events (Text2Event, ), or other n-ary relation models (Text2NKG, ), in the era of LLMs, to preserve richer and more diverse n-ary relations among entities, we propose a new n-ary relation representation $F_{n}=(e_{H},V_{e_{H}})$, utilizing natural language descriptions, instead of structured relations, to represent hyperedges $e_{H}$ among multiple entities $V_{e_{H}}$ as follows.
@@ 5.1 Experimental Setup
Datasets. To evaluate the performance of HyperGraphRAG across multiple domains, we select four knowledge contexts from UltraDomain (UltraDomain, ), as used in LightRAG (LightRAG, ): Agriculture, Computer Science (CS), Legal, and a mixed domain (Mix). In addition, we include the latest international hypertension guidelines (Hypertension, ) as the foundational knowledge for the Medicine domain. For each of the five domains, we sample knowledge fragments one, two, and three hops away to construct questions with ground-truth answers verified by human annotators. We then categorize the questions into Binary Source and N-ary Source, based on whether the sampled knowledge of the question contains facts among $n$ entities ($n>2$). More details can be found in Appendix D.
Baselines. We compare HyperGraphRAG against six publicly available baseline methods: NaiveGeneration (GPT4, ), which directly generates responses using LLM; StandardRAG (RAGsurvey, ), a traditional chunk-based RAG approach; GraphRAG (GraphRAG, ), LightRAG (LightRAG, ), PathRAG (PathRAG, ), and HippoRAG2 (HippoRAG2, ), which are four selected available graph-based RAG methods described in Table 1. To ensure fairness, we use the same generation prompt, which can be found in Appendix E.
Evaluation Metrics. We evaluate the answer accuracy, retrieval efficiency, and generation quality of HyperGraphRAG and its baselines using 3 key metrics: F1, Retrieval Similarity (R-S), and Generation Evaluation (G-E). F1 measures word-level similarity between the generated answer and the ground-truth answer, following FlashRAG (FlashRAG, ). R-S assesses the semantic similarity between the retrieved knowledge and the ground-truth knowledge used to construct the question, in line with RAGAS (RAGAS, ). G-E, inspired by HelloBench (HelloBench, ), is a metric that uses LLM-as-a-judge to evaluate generation quality in 7 dimensions and reports the average score. Details are provided in Appendix E.
Implementation Details. We use OpenAI’s GPT-4o-mini for extraction and generation, and text-embedding-3-small for vector. During retrieval, we set the following parameters: entity retrieval $k_{V}=60$, $\tau_{V}=50$; hyperedge retrieval $k_{H}=60$, $\tau_{H}=5$; and chunk retrieval $k_{C}=5$, $\tau_{C}=0.5$. All experiments were conducted on a server with an 80-core CPU and 512GB RAM.
@@ 5.5 Analysis of Hypergraph Retrieval Efficiency (RQ4)
5.6 Analysis of Hypergraph-Guided Generation Quality (RQ5)
@@ 6 Conclusion
Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant No. 62473271, Grant No. 62176026, and Grant No. 62406036) and the Engineering Research Center of Information Networks, Ministry of Education, China.
@@ B.3 Proof of Proposition 3
$\frac{d}{dL}\mathcal{C}_{H}(L)=\int_{\ell(Y_{H})=L/\eta_{H}}\frac{\partial\mu}{\partial\ell}\cdot\frac{d\ell}{dL}\,d\sigma\geq\int_{\ell(Y_{B})=L/\eta_{B}}\frac{\partial\mu}{\partial\ell}\cdot\frac{d\ell}{dL}\,d\sigma=\frac{d}{dL}\mathcal{C}_{B}(L),$ (38)
@@
$E_{H}(L)-E_{B}(L)=\int_{0}^{L}\frac{d}{d\beta}\left[E_{H}(\beta)-E_{B}(\beta)\right]\,d\beta>0,\quad\text{unless }\mu(\{|F|\geq 3\})=0.$ (42)
@@ Appendix E Evaluation Details
$\text{F1}=\frac{1}{N}\sum_{i=1}^{N}\frac{2\cdot P_{i}\cdot R_{i}}{P_{i}+R_{i}},\quad P_{i}=\frac{|\text{Pred}_{i}\cap\text{GT}_{i}|}{|\text{Pred}_{i}|},\quad R_{i}=\frac{|\text{Pred}_{i}\cap\text{GT}_{i}|}{|\text{GT}_{i}|}$ (43)
@@
$\text{R-S}=\frac{1}{N}\sum_{i=1}^{N}\cos\left(f(k^{(i)}_{\text{retr}}),\;f(k^{(i)}_{\text{gold}})\right)$ (44)
@@
$\text{G-E}=\frac{1}{N}\sum_{i=1}^{N}\operatorname{mean}\left(\frac{1}{7}\sum_{d=1}^{7}s_{i,d},;F1_{i}\right)$ (45)
@@ Appendix G Hyperparameter Settings
For all methods, we adopt a unified set of hyperparameters for all models across both the main evaluation in Table 2 and the time/cost experiments in Table 3 to ensure fair and consistent comparison. For chunk-based methods (e.g., StandardRAG), we retrieve the top-5 chunks using dense similarity. For graph-based methods, including GraphRAG, LightRAG, PathRAG, and HippoRAG2, we retrieve the top-60 relevant elements according to their respective retrieval strategies. HyperGraphRAG performs dual top-60 retrieval over entities and hyperedges, followed by neighborhood expansion. All methods are run using 16 parallel cores and the same generation model (GPT-4o-mini) with temperature 1.0 and a maximum generation length of 32k tokens. Table 4 summarizes the detailed hyperparameter configurations used throughout our experiments.
@@
StandardRAG Chunk 5 chunks 16 GPT-4o-mini